OpenStudy (anonymous):

|-17| + 42 = 40y - |-36| - 39y A. 3 B. 69 C. -69 D. -3

OpenStudy (mertsj):

\[17+42=40y-36-39y\] \[59=y-36\] \[95=y\]

OpenStudy (anonymous):

samething I got..but its not in the answer choics

OpenStudy (agreene):

mertsj is definitely correct--or all 3 of us are wrong somehow.

OpenStudy (anonymous):

95 it is. is there a typo somewhere?

Directrix (directrix):

I also got 95. Check to see if you wrote the problem correctly from the book or wherever.

OpenStudy (mertsj):

Perhaps you made a sign error in the posting.

OpenStudy (bahrom7893):

17 + 42 = 40y - 36 - 39y, merts is right

OpenStudy (anonymous):

NO.....HERE'S THE PIC OF THE QUESTION

OpenStudy (bahrom7893):

4^2 DUUHHH

OpenStudy (agreene):

y=69 4^2 and 42 are rather different.

Directrix (directrix):

I have B) 69 now.

OpenStudy (mertsj):

And how is any sane person supposed to know that 42 is 16?????

Directrix (directrix):

Only the insane know.

OpenStudy (anonymous):

lol you guys are funnny & helpful! You guysrock!!